# Rob A. J. Eyles-Ferris,1 Andrew King,1,2,3 Rhaana L. C. Starling,1 Peter G. Jonker,4 Andrew J. Levan,4 Antonio Martin-Carrillo,5 Tanmoy Laskar,6,4 Jillian C. Rastinejad,7,8 Nikhil Sarin,9,10 Nial R. Tanvir,1 Benjamin P. Gompertz,11,12 Nusrin Habeeb,1 Paul T. O’Brien1 and Massimiliano De Pasquale13

### v1

@@ Abstract
GRB 250702B is a unique astrophysical transient characterised by its nature as a repeating gamma-ray trigger. Its properties include possible periodicity in its gamma-ray light curve, an X-ray counterpart that rose prior to the gamma-ray outbursts and faded quickly, and radio and infrared counterparts. These features are difficult to reconcile with most models of high energy transients but we show that they are compatible with a white dwarf bound to an intermediate mass black hole that is tidally stripped over multiple pericentre passages before being fully disrupted. Accretion onto the black hole powers a mildly relativistic jet that produces the X-rays through internal processes and the infrared and radio counterparts through thermal emission and external shocks respectively but is unable to produce the gamma-ray emission on its own. We propose that chaotic debris streams from the multiple stripping episodes can collide with a period roughly the same as the orbital period of the star. These shocks produce hard X-ray photons that are upscattered by the jet to produce the observed gamma-ray emission. Future analysis of the jet properties will allow us to place firmer constraints on our model.
@@
pubyear: 2025pagerange: Can GRB 250702B be explained as the tidal disruption of a white dwarf by an intermediate mass black hole? Yes.Can GRB 250702B be explained as the tidal disruption of a white dwarf by an intermediate mass black hole? Yes.
@@ 1 Introduction
The localisation offered by EP (hereafter EP) allowed further follow-up of the X-ray counterpart from the Neil Gehrels Swift Observatory (hereafter Swift), EP, NuSTAR and the Chandra X-ray Observatory (hereafter Chandra) (Cheng et al., 2025b; Kennea et al., 2025; O’Connor et al., 2025a, b; Eyles-Ferris et al., 2025). The X-ray counterpart was found to fade rapidly as a power law with temporal index $\alpha\sim 1.9$. A radio counterpart was also identified and has been extensively monitored (Sfaradi et al., 2025a; Bright et al., 2025; Sfaradi et al., 2025b; Atri et al., 2025; Alexander et al., 2025; Tetarenko et al., 2025; Balasubramanian et al., 2025; Grollimund et al., 2025; Rhodes et al., 2025; Levan et al., 2025a).
@@
At optical wavelengths, follow-up was hindered by heavy extinction in the Milky Way and the host. However, Very Large Telescope (VLT) observations were able to identify a fast fading near-infrared (NIR) counterpart (Martin-Carrillo et al., 2025; Levan et al., 2025b, a). The extinction also meant that VLT spectroscopy was insufficient to determine the nature or redshift of the source. Hubble Space Telescope observations, however, conclusively showed the source to be associated with a galaxy. These observations also provided evidence of significant dust lanes which contribute to the extensive extinction (Levan et al., 2025c, a).
@@
Levan et al. (2025a) presented early analysis of GRB 250702B and in particular, showed evidence that the gamma-ray triggers are consistent with period of 2825 s in the observer frame (Levan et al., 2025a). Assuming a relatively low redshift of $z\lesssim 0.3$, they also showed the X-ray, NIR and optical can be attributed to afterglow emission from a GRB or could be linked to a white dwarf (WD) being tidally disrupted and accreted by an intermediate mass black hole (IMBH). These tidal disruption events (TDEs) arise when the WD passes within the tidal radius, $r_{t}=R_{\rm WD}\left(M_{\bullet}/M_{\rm WD}\right)^{1/3}$, of the IMBH. The tidal forces exerted by the IMBH overcome the WD’s self-gravity and it is torn apart (see Maguire et al., 2020, for a review). A significant fraction of the debris remains bound to the IMBH and is accreted to produce a luminous transient. WD-IMBH TDEs have previously been suggested as the progenitor system for ulGRBs (Levan et al., 2014) as the highly super-Eddington accretion rates can offer suitable conditions for a jet to be launched. The fast X-ray transient XRT 000519 (Jonker et al., 2013) may also be an example of such a system and has several remarkable properties and similarities to GRB 250702B, particularly an X-ray precursor and similar evidence of periodicity in its light curve. However, without a redshift for XRT 000519, a direct comparison is difficult.
@@
The host of GRB 250702B has now been observed by the James Webb Space Telescope (JWST, Gompertz et al. in prep.) which has identified the redshift of $z=1.036\pm 0.001$ and allowed the energetics of the transient to be constrained to an isotropic equivalent energy of order $2\times 10^{54}$ erg. However, there is no evidence of transient light and supernovae, in particular, are limited to only the faint outliers of typical GRB-supernovae.
@@
In this work, we reevaluate whether GRB 250702B can indeed be explained by a WD-IMBH TDE armed with more knowledge of the long term evolution and, in particular, the redshift and associated energetics. In Section 2, we analyse X-ray observations from both Swift and Chandra. We present and discuss our model of a repeatedly stripped WD that is finally disrupted to produce GRB 250702B and its counterparts in Section 3 and finally summarise our conclusions in Section 4. Throughout this Letter, we assume a Planck cosmology (Planck Collaboration et al., 2020).
@@ 2 X-ray data analysis
Thanks to the redshift identified by Gompertz et al. (in prep), we are able to place firmer constraints on the X-ray properties and therefore the overall nature of GRB 250702B. Here, we reanalyse the data obtained by Swift-XRT and present our analysis of two Chandra DDT obervations. Errors are given to 90% confidence unless otherwise stated.
@@ 2.1 Swift-XRT
We retrieved the spectral data and count rate light curve of GRB 250702B from the UK Swift Science Data Centre (UKSSDC [^2] https://www.swift.ac.uk/index.php, Evans et al., 2009). These data included all photons detected by Swift-XRT and we fitted them with an absorbed power law model using XSpec v12.13.1 (Arnaud, 1996). We used two absorption components, a Galactic absorber with fixed $N_{H,~{\rm gal}}=3.6\times 10^{21}$ cm-2 (Willingale et al., 2013) and a second absorber at the redshift of GRB 250702B to represent host absorption $N_{H,~{\rm host}}$. We find the spectrum is broadly typical for a transient X-ray source with photon index $\Gamma=1.76\pm 0.21$ and infer a high host column density $N_{H,~{\rm host}}=5.1\pm 1.6\times 10^{22}$ cm-2, consistent with the extensive dust found along the line of sight (Levan et al., 2025a, Gompertz et al. in prep.).
We also checked for spectral evolution by fitting the data from the first epoch of Swift-XRT observations separately. Using the same model, we find $\Gamma=1.56\pm 0.21$ and $N_{H,~{\rm host}}=3.86\pm 1.49\times 10^{22}$ cm-2, consistent with the spectrum reported for EP Follow-up X-ray Telescope (FXT) observations at a consistent time (Cheng et al., 2025b). While these values are consistent within errors of the values found for the whole spectrum, the lower value for $\Gamma$ suggests that the spectrum may have softened somewhat over time. However, due to the low significance of this change, we use the model inferred from our total spectrum fit to convert the count rate light curve to 0.3 - 10 keV luminosities [^3] We also note that the counts to luminosity conversion varies by a factor $\sim 1.01$ between these models and will therefore have negligible impact on our subsequent analysis..
@@ 2.2 Chandra ACIS-S
In addition to Swift, GRB 250702B was also monitored through three Chandra DDT observations of which two are publicly available (O’Connor et al., 2025b; Eyles-Ferris et al., 2025), however, we re-examine these data to account for the now known redshift. Both of these observations were performed using the Advanced CCD Imaging Spectrometer (ACIS-S) instrument. We acquired these data from the Chandra Data Archive and used the Chandra Interactive Analysis of Observations (CIAO) v4.17.0 package (Fruscione et al., 2006) with CALDB v4.12.2 to analyse each observation.
In the case of ObsID 31011, which was taken $\sim 38$ days after the first gamma-ray trigger and lasted for 27.71 ks, the source was detected at a sufficiently high count rate that spectral analysis was feasible. We therefore used the specextract CIAO script to extract the spectrum and the accompanying response files. We fit the same absorbed power law model as for the Swift data but fix $N_{H,~{\rm host}}=5.05\times 10^{22}$ cm-2 due to the small number of counts (29). We find $\Gamma=4.5\pm 1.2$ suggesting the spectrum has continued to soften. Due to the low number of counts and large errors on the inferred parameters, it is difficult to accurately quantify this. In fact, if the spectrum had softened to the extent that we measure with this fit, it suggests that GRB 250702B’s X-ray counterpart had transitioned to a likely thermal phase, rebrightened compared to the last Swift-XRT epoch and plateaued. This is inconsistent with our analysis of ObsID 31468. We therefore assume the same spectrum as the Swift-XRT data to convert from our measured count rate of $9.75\pm 1.94\times 10^{-4}$ count s-1 to the 0.3-10 keV luminosity, similarly to O’Connor et al. (2025b).
By the time Chandra observed ObsID 31468 at  65 days post trigger, GRB 250702B’s counterpart had faded significantly. Instead of attempting to extract a spectrum, we therefore used the srcflux CIAO script to measure a count rate of $1.12^{+1.27}_{-0.74}\times 10^{-4}$ count s-1. Again, we use the same spectrum as the Swift-XRT observations to convert this to a 0.3-10 keV luminosity. We find that the observed luminosity is somewhat lower than expected from extrapolating the X-ray light curve. However, this is of low significance ($\sim 2.5$-$\sigma$) and could be a Poisson effect due to the small number of counts.
We show the complete X-ray light curve, including the data from EP observations described by Cheng et al. (2025a) and Cheng et al. (2025b), in the context of a WD-IMBH TDE in Figure 1.
@@ 3 A WD-IMBH TDE model of GRB 250702B
We now examine GRB 250702B further in the context of a WD being disrupted by an IMBH and whether these features can be produced by such an event. We make several key assumptions - the periodicity investigated by Levan et al. (2025a) is both real and related to the orbital period of the WD, and that the conditions, such as the properties of the magnetic field near the IMBH, are sufficient to launch a relativistic jet. Importantly, as shown by Levan et al. (2025a), the orbital period eliminates both supermassive black holes and main sequence stars as components of the progenitor system.
@@ 3.1 Repeated tidal stripping and accretion of a white dwarf
A key property of WDs is the inverted relationship between their mass and their radius - lower mass WDs have larger radii and vice versa (e.g. Ambrosino, 2020). This particular property offers an explanation for the precursor X-ray emission observed in GRB 250702B. Analysis of extreme mass ratio inspirals (EMRIs, Amaro-Seoane, 2018) show that while gravitational wave emission does result in a decay in the orbital semi-major axis, this effect is slow and therefore the WD’s orbit is, at least initially, relatively stable (e.g. Peters, 1964).
@@
However, as the WD undergoes a pericentre passage, while its core may be outside the tidal radius $r_{t}$, it is possible that the outer layers may be sufficiently weakly bound that a small amount of mass is stripped and accreted by the IMBH. There is also a possibility of thermonuclear ignition as the WD is compressed by the tidal force of the IMBH (e.g. Rosswog et al., 2008, 2009; MacLeod et al., 2016) although this effect is likely small for the shallow encounter assumed here. Assuming the WD is able to expand quickly enough to quench any thermonuclear burning, i.e. $t_{\rm dyn}<t_{\rm burning}$ (Rosswog et al., 2009), the now lower mass WD will therefore expand in radius, with some impact from tidal distortion also possible (e.g. Maguire et al., 2020). This stripping will be repeated on the subsequent pericentre passages and is a runaway effect - the WD will have an incrementally larger radius on each passage and therefore more matter will be stripped - until the WD has lost sufficient mass that it will lie within $r_{t}$ and be fully disrupted. To illustrate this, we show the evolution of $r_{t}$ with WD and IMBH mass in Figure 2. This model is similar to that proposed by King (2022, see also () and ()) to explain quasi-periodic eruptions (QPEs) from galactic nuclei. In fact, it is likely that the early tidal stripping events may resemble QPEs as the accretion rate will decay significantly quicker than a later full disruption. King (2022) shows that for most systems, the mass transfer is stable and leads to an expanding orbit to conserve angular momentum. However, in this case, the orbital period is significantly shorter than the viscous timescale (of order several thousand seconds) which means that an accretion disk will not be efficiently formed before the WD returns to pericentre. In this case, then, angular momentum cannot be transferred back to the WD and the orbital semi-major axis does not expand. It is important to also note that although the stripping episodes result in fallback rates that evolve quicker than that from the final disruption, they can achieve comparable peak accretion rates (Nixon et al., 2021).
Figure 2: The evolution of the tidal radius with IMBH and WD mass derived from the relations between WD mass and radius (e.g. Ambrosino, 2020). Note the WD will move left along these curves with repeated tidal stripping assuming $t_{\rm dyn}<t_{\rm burning}$.
@@
Due to the repeated stripping events, shocks between the different debris streams are still likely to induce the formation of an accretion flow at early times (Rosswog et al., 2009; Bonnerot & Stone, 2021), although the transfer of angular momentum from a particular debris stream still cannot occur before the next pericentric passage and the runaway effect still occurs. This means that material from later passages will eventually be accreted and, due to the low mass of the IMBH involved, will result in highly super-Eddington accretion. In fact, assuming a $10^{4}~M_{\odot}$ IMBH, the Eddington luminosity is of order $10^{42}$ erg s-1. The vast majority of the stripped material will be accreted over $\lesssim 10^{6}$ s so stripping and accreting as little $10^{-5}~M_{\odot}$ will result in super-Eddington accretion. The timescale over which the jet formation occurs is unclear and due to the complex nature of the WD’s radius evolution and the amount of matter stripped, we leave this to a future study. However, a jet may only be launched only a few tens of passages prior to the full disruption of the WD as disordered magnetic fields may require the accretion rate to be several orders of magnitude super-Eddington.
@@
The orbit of the WD will naturally differ from that expected in other TDE systems. Typically a parabolic ($e=1$) orbit is assumed but this results in a single encounter as the star is disrupted. Instead, we require the WD to be bound to the IMBH to ensure continuous orbits (i.e. $e<1$). Achieving this likely requires Hills capture where the WD is initially in a binary system that is disrupted by the IMBH leaving only the WD bound (Hills, 1988) and has been suggested in the case of other repeating transients (e.g. King, 2022; Cufari et al., 2022). If we assume a $0.6~M_{\odot}$ WD, close to the peak of the WD mass distribution, orbiting an IMBH with mass $10^{4}~M_{\odot}$ and that $r_{p}=r_{t}$, for the inferred period, we require $e\simeq 0.945$. As typically the more massive star is captured, we require either a WD binary or a WD in a binary with a low mass main sequence star such as an M dwarf. The fact that a presumed runaway is observed in GRB 250702B implies that the WD’s initial orbit was very close to its final orbit and it likely was bound to the IMBH only shortly before its disruption. We examine the likelihood of achieving this orbit further in Section 3.4.

### v2

@@ Abstract
GRB 250702B is a unique astrophysical transient characterised by its nature as a repeating gamma-ray trigger. Its properties include possible periodicity in its gamma-ray light curve, an X-ray counterpart that rose prior to the gamma-ray outbursts and faded quickly, and radio and infrared counterparts. These features are difficult to reconcile with most models of high energy transients but we show that they are compatible with a white dwarf bound to an intermediate mass black hole that is tidally stripped over multiple pericentre passages before being fully disrupted. In this model, accretion onto the black hole powers a mildly relativistic jet that produces the X-rays through internal processes and the infrared and radio counterparts through thermal emission and external shocks respectively but is unable to produce the gamma-ray emission on its own. We find that chaotic debris streams from the multiple stripping episodes can collide with a period roughly the same as the orbital period of the star. These shocks produce X-ray photons that are upscattered by the jet to produce the observed MeV gamma-ray emission. Future analysis of the jet properties will allow us to place firmer constraints on our model.
@@
pubyear: 2025pagerange: Can GRB 250702B be explained as the tidal disruption of a white dwarf by an intermediate mass black hole? Yes.Data Availability
@@ 1 Introduction
The localisation offered by EP allowed further follow-up of the X-ray counterpart from the Neil Gehrels Swift Observatory (hereafter Swift), EP, NuSTAR and the Chandra X-ray Observatory (hereafter Chandra) (GCN40917; GCN40919; GCN41014; GCN41309; GCN41767). The X-ray counterpart was found to fade rapidly as a power law with temporal index $\alpha\sim 1.9$. A radio counterpart was also identified and has been extensively monitored (GCN40979; GCN40985; GCN41053; GCN41054; GCN41059; GCN41061; GCN41145; GCN41147; GCN41215; Levan25).
@@
At optical wavelengths, follow-up was hindered by significant extinction in the Milky Way and the host. However, Very Large Telescope (VLT) observations were able to identify a fast fading near-infrared (NIR) counterpart (GCN40924; GCN40961; Levan25). Hubble Space Telescope observations conclusively showed the source to be associated with a galaxy. These observations also provided evidence of significant dust lanes which contribute to the extensive extinction (GCN41096; Levan25).
@@
Levan25 presented early analysis of GRB 250702B and in particular, showed evidence that the multiple gamma-ray triggers are consistent with a period of 2825 s in the observer frame (Levan25). Assuming a relatively low redshift of $z\lesssim 0.3$, they also showed the X-ray, NIR and optical can be attributed to afterglow emission from a GRB or could be linked to a white dwarf (WD) being tidally disrupted and accreted by an intermediate mass black hole (IMBH). These tidal disruption events (TDEs) arise when the WD passes within the tidal radius, $r_{t}=R_{\rm WD}\left(M_{\bullet}/M_{\rm WD}\right)^{1/3}$, of the IMBH. The tidal forces exerted by the IMBH overcome the WD’s self-gravity and it is torn apart (see Maguire20, for a review). A significant fraction of the debris remains bound to the IMBH and is accreted to produce a luminous transient. WD-IMBH TDEs have previously been suggested as the progenitor system for ulGRBs (Levan14) as the highly super-Eddington accretion rates can offer suitable conditions for a jet to be launched. The fast X-ray transient XRT 000519 (Jonker13) may also be an example of such a system and has several remarkable properties and similarities to GRB 250702B, particularly an X-ray precursor and similar evidence of periodicity in its light curve. However, without a redshift for XRT 000519, a direct comparison is difficult.
@@
The host of GRB 250702B has now been observed by the James Webb Space Telescope (JWST, Gompertz25) which has identified the redshift of $z=1.036\pm 0.004$ and allowed the energetics of the transient to be constrained to an isotropic equivalent energy of order $2\times 10^{54}$ erg. However, there is no evidence of transient light and supernovae, in particular, are limited to only the faint outliers of typical GRB-supernovae. The availability of the redshift has allowed several models of GRB 250702B’s behaviour to be assessed although its nature remains ambiguous (see e.g. Carney25; OConnor25, for an overview). These include a micro-TDE where a main sequence star is disrupted by a stellar mass black hole (Beniamini25; OConnor25) and a merger between a stellar mass black hole (Neights25) while a WD-IMBH TDE remains a possibility (Li25).
@@
In this work, we reevaluate whether GRB 250702B can indeed be explained by a WD-IMBH TDE armed with more knowledge of the long term evolution and, in particular, the redshift and associated energetics. A similar model is examined in Li25, although the assumptions made there differ significantly in some areas as briefly discussed in Section 3.1. In Section 2, we analyse X-ray observations from both Swift and Chandra. We present and discuss our model of a repeatedly stripped WD that is finally disrupted to produce GRB 250702B and its counterparts in Section 3 and finally summarise our conclusions in Section 4. Throughout this Letter, we assume a Planck cosmology (Planck20).
@@ 2 X-ray data analysis
Thanks to the redshift identified by Gompertz25, we are able to place firmer constraints on the X-ray properties and therefore the overall nature of GRB 250702B. Here, we reanalyse the data obtained by Swift-XRT and present our analysis of two Chandra DDT obervations. Errors are given to 90% confidence unless otherwise stated.
@@ 2.1 Swift-XRT
We retrieved the spectral data and count rate light curve of GRB 250702B from the UK Swift Science Data Centre (UKSSDC [^2] https://www.swift.ac.uk/index.php, Evans09). These data included all photons detected by Swift-XRT and we fitted them with an absorbed power law model using XSpec v12.13.1 (Arnaud96). We used two absorption components, a Galactic absorber with fixed $N_{H,~{\rm gal}}=3.6\times 10^{21}$ cm-2 (Willingale13) and a second absorber at the redshift of GRB 250702B to represent host absorption $N_{H,~{\rm host}}$. We find the spectrum is broadly typical for a transient X-ray source with photon index $\Gamma=1.76\pm 0.21$ and infer a high host column density $N_{H,~{\rm host}}=5.1\pm 1.6\times 10^{22}$ cm-2, consistent with the extensive dust found along the line of sight (Levan25; Gompertz25).
We also checked for spectral evolution by fitting the data from the first epoch of Swift-XRT observations separately. Using the same model, we find $\Gamma=1.56\pm 0.21$ and $N_{H,~{\rm host}}=3.86\pm 1.49\times 10^{22}$ cm-2, consistent with the spectrum reported for EP Follow-up X-ray Telescope (FXT) observations at a contemporary time (GCN40917). We therefore find no significant evidence for spectral evolution and use the model inferred from our total spectrum fit to convert the count rate light curve to 0.3 - 10 keV luminosities [^3] We also note that the counts to luminosity conversion varies by a factor $\sim 1.01$ between these models and will therefore have negligible impact on our subsequent analysis..
@@ 2.2 Chandra ACIS-S
In addition to Swift, GRB 250702B was also monitored through three Chandra DDT observations of which two are publicly available (GCN41309; GCN41767), however, we re-examine these data to account for the now known redshift. Both of these observations were performed using the Advanced CCD Imaging Spectrometer (ACIS-S) instrument. We acquired these data from the Chandra Data Archive and used the Chandra Interactive Analysis of Observations (CIAO) v4.17.0 package (Fruscione06) with CALDB v4.12.2 to analyse each observation.
ObsID 31011 was taken $\sim 38$ days after the first gamma-ray trigger and lasted for 27.71 ks. We used the specextract CIAO script to extract the spectrum and the accompanying response files but due to a gain calibration issue (private communication), we do not use these for spectral analysis. Instead, we assume the same spectrum as the Swift-XRT data to convert from our measured count rate of $9.75\pm 1.94\times 10^{-4}$ count s-1 to the 0.3-10 keV luminosity, similarly to GCN41309.
By the time Chandra observed ObsID 31468 at  65 days post trigger, GRB 250702B’s counterpart had faded significantly. Instead of attempting to extract a spectrum, we therefore used the srcflux CIAO script to measure a count rate of $1.12^{+1.27}_{-0.74}\times 10^{-4}$ count s-1. Again, we use the same spectrum as the Swift-XRT observations to convert this to a 0.3-10 keV luminosity.
We show the complete X-ray light curve, including the data from EP observations described by Li25, in the context of a WD-IMBH TDE in Figure 1.
@@ 3 A WD-IMBH TDE model of GRB 250702B
We now examine GRB 250702B further in the context of a WD being disrupted by an IMBH and whether these features can be produced by such an event. We make several key assumptions - the periodicity investigated by Levan25 is both real and related to the orbital period of the WD [^4] This implies a significantly shorter orbital period than assumed in Li25., and that the conditions, such as the properties of the magnetic field near the IMBH, are sufficient to launch a relativistic jet. Importantly, as shown by Levan25, the orbital period eliminates both supermassive black holes and main sequence stars as components of the progenitor system.
@@ 3.1 Repeated tidal stripping and accretion of a white dwarf
A key property of WDs is the inverted relationship between their mass and their radius - lower mass WDs have larger radii and vice versa (e.g. Ambrosino20). This particular property offers an explanation for the precursor X-ray emission observed in GRB 250702B. Analysis of extreme mass ratio inspirals (EMRIs, AmaroSeoane18) show that while gravitational wave emission does result in a decay in the orbital semi-major axis, this effect is slow and therefore the WD’s orbit is, at least initially, relatively stable (e.g. Peters64).
@@
However, as the WD undergoes a pericentre passage, while its core may be outside the tidal radius $r_{t}$, it is possible that the outer layers may be sufficiently weakly bound that a small amount of mass is stripped and accreted by the IMBH. Assuming the WD is able to expand quickly enough to quench any thermonuclear burning which can be induced through compression of the WD in deep encounters (Rosswog08; Rosswog09; Macleod16), i.e. $t_{\rm dyn}<t_{\rm burning}$, the now lower mass WD will therefore expand in radius. There may also be additional expansion due to tidal distortion although fully quantifying this effect is beyond the scope of this work (e.g. Maguire20). This stripping will be repeated on the subsequent pericentre passages and is a runaway effect - the WD will have an incrementally larger radius on each passage and therefore more matter will be stripped - until the WD has lost sufficient mass that it will lie within $r_{t}$ and be fully disrupted. This model is similar to that proposed by King22 to explain quasi-periodic eruptions (QPEs) from galactic nuclei. In fact, it is likely that the early tidal stripping events may resemble QPEs as the accretion rate will decay significantly quicker than a later full disruption.
King22 shows that for most systems, the mass transfer is stable and leads to an expanding orbit to conserve angular momentum. However, in this case, the orbital period is significantly shorter than the viscous timescale (from Equation 25 of King22 and assuming a modest $H/R\sim 0.1$, we find it to be of order several thousand seconds, significantly shorter than that assumed by Li25) which means that an accretion disk will not be efficiently formed before the WD returns to pericentre. In this case, then, angular momentum cannot be transferred back to the WD and the orbital semi-major axis does not expand. The short orbital period is crucial in this case and is the direct cause of the runaway. It is important to also note that although the stripping episodes result in fallback rates that evolve quicker than that from the final disruption, they can achieve comparable peak accretion rates (Nixon21).
@@
Due to the repeated stripping events, shocks between the different debris streams are still likely to induce the formation of an accretion flow at early times (Rosswog09; Bonnerot21b) [^5] However, the transfer of angular momentum from a particular debris stream still cannot occur before the next pericentric passage due to the long viscous timescale compared to the orbital period. This means the WD will maintain its orbit and the runaway stripping episodes will continue. This means that material from later passages will eventually be accreted and, due to the low mass of the IMBH involved, will result in highly super-Eddington accretion. In fact, assuming a $10^{4}~M_{\odot}$ IMBH, the Eddington luminosity is of order $10^{42}$ erg s-1. From Figure 1, the vast majority of the stripped material will be accreted over $\lesssim 10^{6}$ s so stripping and accreting as little $10^{-5}~M_{\odot}$ will result in super-Eddington accretion. The timescale over which the jet launch, and therefore bright X-ray emission becomes observable, occurs is unclear and due to the complex nature of the WD’s radius evolution and the amount of matter stripped, we leave this to a future study. However, the jet may only be launched only a few tens of passages prior to the full disruption of the WD as disordered magnetic fields may require the accretion rate to be several orders of magnitude super-Eddington.
@@
The orbit of the WD will naturally differ from that expected in other TDE systems. Typically a parabolic ($e=1$) orbit is assumed but this results in a single encounter as the star is disrupted. Instead, we require the WD to be bound to the IMBH to ensure continuous orbits (i.e. $e<1$). Achieving this likely requires Hills capture where the WD is initially in a binary system that is disrupted by the IMBH leaving only the WD bound (Hills88) and has been suggested in the case of other repeating transients (e.g. King22; Cufari22). We discuss this further in Section 3.4.